OpenStudy (anonymous):

Jim can fill a pool carrying buckets of water in 30 minutes. Sue can do the same job in 45 minutes. Tony can do the same job in 1 ½ hours. How quickly can all three fill the pool together?

OpenStudy (anonymous):

For this you can create a chart

OpenStudy (bahrom7893):

lol everyone feelin too lazy to start?

OpenStudy (anonymous):

20 minutes

OpenStudy (anonymous):

no wait.

OpenStudy (anonymous):

45 minutes

OpenStudy (mertsj):

\[\frac{x}{.5}+\frac{x}{.75}+\frac{x}{1.5}=1\]

OpenStudy (mertsj):

Solve that.

OpenStudy (mertsj):

\[x=\frac{1}{4} hr\]

OpenStudy (anonymous):

thanks for your help :D

OpenStudy (mertsj):

yw